OpenStudy (anonymous):

Rewrite with only sin x and cos x. sin 3x - cos x

OpenStudy (tkhunny):

Have you considered using \(\sin(a+b) = \sin(a)\cos(b) + \cos(a)\sin(b)\)?

OpenStudy (anonymous):

Yes I have but I am still like.. Unsure as to how to get the answer

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

Would it just be sin(3x-x) which is sin(2x)? @tkhunny

OpenStudy (anonymous):

woops. sorry. i meant sin(3x+x) which is sin (4x) @tkhunny

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@ganeshie8 @mathstudent55

OpenStudy (tkhunny):

\(\sin(3x) = \sin(x + 2x)\)